OpenStudy (jdanmath):

Find the limit if it exists lim as x-> infinity 4/(x-3)^3

OpenStudy (johnweldon1993):

What do you think the answer is?

OpenStudy (jdanmath):

I already know the answer- 0, but I don't understand how. Limits at infinity get me lost.

OpenStudy (johnweldon1993):

Okay think about it like this Since we want to know the limit as we approach infinity...aka REALLY BIG NUMBERS!!! lets say I plug in 1000 for 'x' \[\large \frac{4}{(1000 - 3)^3}\] That will come out to a REALLY small number right? What about if I plug in 10000 now? \[\large \frac{4}{(10000 - 3)^3}\] That would get even smaller still right? As values of 'x' get larger and larger approaching infinity....the limit gets smaller and smaller approaching zero

OpenStudy (jdanmath):

okay i understand this but now im confused between limits at infinity and infinite limits :(

OpenStudy (johnweldon1993):

Not sure what you mean, do you have an example?

OpenStudy (jdanmath):

limit as x approaches infinity vs limit as x approaches a number(from + or -)

OpenStudy (jdanmath):

limas x->-1+ 2/(x+1) answer is infinity

OpenStudy (jdanmath):

and in the previous answer i asked you the answer was 0

OpenStudy (johnweldon1993):

Sorry for the extremely late response! Lets work with that example you gave \[\large lim_{x \rightarrow -1^+} \frac{2}{(x + 1)}\] What would happen if you plug in -1 into that equation for 'x'? \[\large \frac{2}{(-1 + 1)} = \frac{-2}{0}\] What do we know about dividing by 0? The answer is infinite Why? think about it...if I plug lets say 1 in for x in the equation \(\large \frac{1}{x}\) the answer is 1...what about if I plug in 1/2? the answer would be 2 What about 1/4? the answer would be 4 What about 1/1000? The answer would be 1000 etc... as the numbers get smaller and smaller *closer to 0* the value of the expression gets bigger and bigger *approaching infinity*

OpenStudy (jdanmath):

Thank you so much! You're awesome :)

OpenStudy (polly7000):

It's zero. Just a starting tip. You should consider both the situations when x approaches positive and negative infinity.